# Polyp-Gen: Realistic and Diverse Polyp Image Generation for Endoscopic Dataset Expansion

### v1

@@ I INTRODUCTION
Early detection of polyps in the lower gastrointestinal tract (GI) can significantly reduce the risk of developing life-threatening colorectal cancer. During this procedure, gastroenterologists typically need to identify polyps from lengthy videos, which is greatly reliant on their expertise and requires considerable labor. With the development of artificial intelligence and deep learning techniques, automated diagnostic systems (ADS) demonstrate significant potential in endoscopic examinations [1, 2, 3]. The efficacy of ADS is highly dependent on the quantity and quality of the training data. However, due to high manual annotation costs and strict privacy protection requirements, acquiring high-quality data with realism and diversity poses a considerable challenge within the medical domain.
@@
tations, we introduce Polyp-Gen, the first full-automatic diffusion-based framework to create a wide range of realistic endoscopic images for data augmentationsetexpansion. Specifically, we propose a spatial-aware diffusion training scheme to enhance the structural context of polyp boundary regions while preserving endoscopic global information, enabling the transformation between polyp images and normal images, as shown in Fig. 1. In addition, to capture medical priors for the automatic identification of potential polyp locations, we devise a hierarchical retrieval-based sampling strategy with a mask proposer to match similar fine-grained spatial features. In this way, our generated images can achieve quality comparable to real images and enhance downstream diagnosis tasks. The contributions of our work can be summarised as follows:
@@ II RELATED WORK
GAN-based models. Early studies [4, 7, 8, 9] adopted Generative Adversarial Networks (GAN) [12] for endoscopic image generation, especially polyp image generation. Among them, SinGAN-Seg [7] was trained using a single polyp image and corresponding mask to produce synthetic samples. However, this is a time-intensive procedure, and the resulting images closely resemble the original image. SemanticPolypGAN [8] utilized masked images as conditional input to generate polyp images, and Hemin et al. [9] achieved a conversion between polyp image and normal image by a simple U-Net based conditional GAN. PolypConnect [4] devised a polyp inpainting pipeline by extracting and combining edge information.
@@ III-B Stable diffusion model
We implement our Polyp-Gen based on the Stable Diffusion (SD) [15], which demonstrates superior ability in text-guided image generation tasks. The SD is a two-stage diffusion model that contains a variational autoencoder (VAE) and a U-Net denoiser. The forward and backward process of SD is performed in the latent space allowing for better control and manipulation of the generated content. Given an input image $I$, the VAE encoder extracts the latent representation of the image $z=E(I)\in\mathcal{R}^{\frac{h}{8}\times\frac{w}{8}\times 4}$. In the forward process, a noisy latent variable $z_{t}$ is obtained by introducing random noise $\epsilon\sim N(0,1)$ at time step $t$, given by $z_{t}=\alpha_{t}z_{t-1}+\sigma_{t}\epsilon$, where $\alpha_{t}$ and $\sigma_{t}$ are coefficients that govern the noise schedule. Subsequently, the conditional diffusion model $\epsilon_{\theta}$ is optimized using the following objective:
$L_{D}=\mathbb{E}_{z,c,\epsilon\sim N(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(z_{t},c,t\right)\right\|_{2}^{2}\right],$ (1)
@@ III-C Spatial-aware diffusion training
$L_{MSE}=\mathbb{E}_{z,c,\epsilon\sim N(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\left[z_{t},z^{m},m\right],c,t\right)\right\|_{2}^{2}\right].$ (2)
@@
$L_{LG}=\mathbb{E}_{z,c,\epsilon\sim N(0,1),t}\left[\left\|m\odot\epsilon-m\odot\epsilon_{\theta}\left(\left[z_{t},z^{m},m\right],c,t\right)\right\|_{2}^{2}\right].$ (3)
@@ III-D Hierarchical retrieval-based sampling
=$\mathcal{M}=\{p:p=\mathop{\arg\min}\limits_{u}d(u,v),I^{v}_{c}\in M_{P}\odot I_{c}\},$ (5)
@@
where $j,k\in\{1,2,\cdots,N\}$, $N=\frac{H}{P}\times\frac{W}{P}$ denotes the number of patches, and $P$ is the patch size. The $p$-th patch corresponds to the $(h,w)$ coordinates in the query image space. To further obtain similar local structures from $\mathcal{M}$, we employ the density-based spatial clustering of applications with noise (DBSCAN) [24] algorithm to cluster points that exhibit high proximity. The radius $\epsilon_{r}$ is set as $12\times P+$ 1. We then select the cluster with the maximum number of points and obtain its enclosing rectangle as the mask proposal $M_{q}$. The position of the mask, representing the most similar local structure to the potential polyp region in the reference image, also indicates the most plausible location for polyp generation in the query image. After obtaining the mask proposal $M_{q}$, the trained Polyp-Gen can generate polyp images $I^{\prime}_{P}=\theta(I_{q},M_{q},\textit{Polyp})$.
@@ IV-C Downstream task evaluation
The generated polyp images can be directly employed in training polyp detection models without requiring mask annotations. Additionally, a pixel-wise mask can be acquired through semi-supervised segmentation models like Segment Anything (SAM) [35]. This is crucial as it reduces the manual labeling costs associated with expert annotation of masks in medical data.

### v3

@@ I INTRODUCTION
Early detection of polyps in the lower gastrointestinal tract (GI) can significantly reduce the risk of developing life-threatening colorectal cancer. During this procedure, gastroenterologists typically need to identify polyps from lengthy videos, which is greatly reliant on their expertise and requires considerable labor. With the development of artificial intelligence and deep learning techniques, automated diagnostic systems (ADS) demonstrate significant potential in endoscopic examinations [1, 2, 3, 4]. The efficacy of ADS is highly dependent on the quantity and quality of the training data [5, 6]. However, due to high manual annotation costs and strict privacy protection requirements, acquiring high-quality data with realism and diversity poses a considerable challenge within the medical domain.
@@
To address these limitations, we introduce Polyp-Gen, the first full-automatic diffusion-based framework to create a wide range of realistic endoscopic images for dataset expansion. Specifically, we propose a spatial-aware diffusion training scheme to enhance the structural context of polyp boundary regions while preserving endoscopic global information, enabling the transformation between polyp images and normal images, as shown in Fig. 1. In addition, to capture medical priors for the automatic identification of potential polyp locations, we devise a hierarchical retrieval-based sampling strategy with a mask proposer to match similar fine-grained spatial features. In this way, our generated images can achieve quality comparable to real images and enhance downstream diagnosis tasks. The contributions of our work can be summarised as follows:
@@ II RELATED WORK
GAN-based models. Early studies [7, 10, 11, 12] adopted Generative Adversarial Networks (GAN) [15, 16, 17, 18] for endoscopic image generation, especially polyp image generation. Among them, SinGAN-Seg [10] was trained using a single polyp image and corresponding mask to produce synthetic samples. However, this is a time-intensive procedure, and the resulting images closely resemble the original image. SemanticPolypGAN [11] utilized masked images as conditional input to generate polyp images, and Hemin et al. [12] achieved a conversion between polyp image and normal image by a simple U-Net based conditional GAN. PolypConnect [7] devised a polyp inpainting pipeline by extracting and combining edge information.
@@ III-B Stable diffusion model
We implement our Polyp-Gen based on the Stable Diffusion (SD) [21], which demonstrates superior ability in text-guided image generation tasks. The SD is a two-stage diffusion model that contains a variational autoencoder (VAE) and a U-Net denoiser. The forward and backward process of SD is performed in the latent space allowing for better control and manipulation of the generated content. Given an input image II, the VAE encoder extracts the latent representation of the image $z=E(I)\in\mathcal{R}^{\frac{h}{8}\times\frac{w}{8}\times 4}$. In the forward process, a noisy latent variable $z_{t}$ is obtained by introducing random noise $\epsilon\sim N(0,1)$ at time step $t$, given by $z_{t}=\alpha_{t}z_{t-1}+\sigma_{t}\epsilon$, where $\alpha_{t}$ and $\sigma_{t}$ are coefficients that govern the noise schedule. Subsequently, the conditional diffusion model $\epsilon_{\theta}$ is optimized using the following objective:
$L_{D}=\mathbb{E}_{z,c,\epsilon\sim N(0,1),t}\left\|\epsilon-\epsilon_{\theta}\left(z_{t},c,t\right)\right\|_{2}^{2},$ (1)
@@ III-C Spatial-aware diffusion training
$L_{MSE}=\mathbb{E}_{z,c,\epsilon\sim N(0,1),t}\left\|\epsilon-\epsilon_{\theta}\left(\left[z_{t},z^{m},m\right],c,t\right)\right\|_{2}^{2},$ (2)
@@
$L_{LG}=\mathbb{E}_{z,c,\epsilon\sim N(0,1),t}\left\|m\odot\epsilon-m\odot\epsilon_{\theta}\left(\left[z_{t},z^{m},m\right],c,t\right)\right\|_{2}^{2}$ (3)
@@ III-D Hierarchical retrieval-based sampling
$\mathcal{M}=\{p:p=\mathop{\arg\min}\limits_{u}d(u,v),I^{v}_{c}\in M_{P}\odot I_{c}\},$ (5)
@@
where $j,k\in\{1,2,\cdots,N\}$, $N=\frac{H}{P}\times\frac{W}{P}$ denotes the number of patches, and $P$ is the patch size. The $p$-th patch corresponds to the $(h,w)$ coordinates in the query image space. To further obtain similar local structures from $\mathcal{M}$, we employ the density-based spatial clustering of applications with noise (DBSCAN) [30] algorithm to cluster points that exhibit high proximity. The radius $\epsilon_{r}$ is set as $2\times P+1$. We then select the cluster with the maximum number of points and obtain its enclosing rectangle as the mask proposal $M_{q}$. The position of the mask, representing the most similar local structure to the potential polyp region in the reference image, also indicates the most plausible location for polyp generation in the query image. After obtaining the mask proposal $M_{q}$, the trained Polyp-Gen can generate polyp images $I^{\prime}_{P}=\theta(I_{q},M_{q},\textit{Polyp})$.
@@ IV-C Downstream task evaluation
The generated polyp images can be directly employed in training polyp detection models without requiring mask annotations. Additionally, a pixel-wise mask can be acquired through semi-supervised segmentation models like Segment Anything (SAM) [41, 42, 43]. This is crucial as it reduces the manual labeling costs associated with expert annotation of masks in medical data.
@@ V CONCLUSION
VI Acknowledgment
This work was supported by Innovation and Technology Commission - Innovation and Technology Fund ITS/229/22 and Hong Kong Research Grants Council (RGC) General Research Fund 14220622.